HUBBLE SPACE TELESCOPE DAILY REPORT #1969 PERIOD COVERED: 0000Z (UTC) 09/23/97 - 0000Z (UTC) 09/24/97 Daily Status Report as of 267/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD 7659 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make dark calibration images in order to update the acquisition bad pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 7792 (HST Support for Mars Global Surveyor) The WF/PC-2 was used to make imaging observations of the planet Mars in support of the Mars Global Surveyor Mission during the aero-braking by monitoring dust opacity on the planet's atmosphere. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD 7783 (STIS) The Space Telescope Imaging Spectrograph (CCD) was used to make observations taken in parallel with the prime science. The observations were completed, and no problems were reported. 1.5 Completed NIC/1 7693 (NICMOS Pupil Transfer Function) The Near Infrared Camera (Camera One) was used to make observations of P330E-Cal in order to perform pupil transfer function observations. The output of this program will be a two-dimensional surface fit that will give the large scale corrections needed to flat-field the images to an accuracy of about 2%. Also, changes in the PSF will be measured as a function of position within the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxies 3C136.1 and 3C133.0 . Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena; they may be used as probes of their environments; and they are unique probes of the early universe. The observations were completed as planned, and no problems were reported. 1.7 Completed STIS/MA1 7091 (STIS Pixel-To-Pixel Response Stability) The Space Telescope Imaging Spectrograph (MA1) was used to make observations of standard star BD8D4211 in order to determine the pixel-to-pixel response stability. This is to illustrate that STIS has the capability of achieving signal to noise ratio of 100:1 in both Band 1 {115-170 nm} and Band 2 {165 - 310 nm}. The observations were completed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of NIC/2/3 7791 (NICMOS Focus Sweep at 2 FOM positions) The Near Infrared Camera (Cameras Two and Three) was used to make observations of the cluster called Target-2 as part of a NICMOS focus sweep. The main objectives of this proposal are: a} to cover sufficient region of PAM focus space to ensure the optimal locations are included, and b} measure vignetting and image quality at different FOM positions. The observations were executed as schedule, and no anomalies were noted. 1.9 Completed Two Sets of NIC/1/2/3 7729 (NICMOS Short One Orbit Parallel) The Near Infrared Camera (Cameras One, Two, and Three) was used to obtain "JHK" imaging of sky regions while another instrument was prime. These observations were part of the short NICMOS one-orbit parallel program. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Three Sets of WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.11 Completed STIS/MA2 7604 (Cycle 7 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make MAMA dark measurements. These darks will be used in later image reductions and also trended over the instrument's life time as an indicator of the detector's degradation or damage. The observations were completed as scheduled, and no problems were reported. 1.12 Completed NIC/2 7420 (A Near-Infrared Search for Very Low Mass Companions to Stars Within 10 Parsecs of the Sun) The Near Infrared Camera (Camera Two) was used to make searching observations for very low mass companions to variable star BD+25D3173. This star is known to be within 10 parsecs of the Sun. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed NIC/2 7418 (Imaging of Younf Stellar Object Circumstellar Nebulosity) WF/PC-2 observations have shown that several young stellar objects in the nearby Taurus molecular clouds are seen only indirectly via scattered light. New NICMOS (Camera Two) images will be compared with the existing WFPC2 images of IRAS04248 and with model reflection nebulae to constrain the circumstellar density distributions. The dark absorption lines already seen in these systems suggest that the NICMOS images will reveal edge-on disks similar to the HST results for HH 30, and thus allow direct observation of the disk geometry. The observations were completed with no reported anomalies 1.14 Completed Three Sets of WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: A TTR was written for an occurrence at 267/0208Z when no UPDs or GCMRs were possible due to an NCC problem. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH